OpenStudy (anonymous):

how to prove the identity of 1-cos(x)/sin(x) + sin(x)/1+cos(x) = 2 csc (x)

OpenStudy (zzr0ck3r):

\[\frac{1-cos(x)}{sin(x)}+\frac{sin(x)}{1+cos(x)}=2csc(x)\]

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

@cgurley86

OpenStudy (anonymous):

yeah thats the equation

OpenStudy (zzr0ck3r):

its not

OpenStudy (zzr0ck3r):

are you sure its 1+cos(x) in the second term?

OpenStudy (anonymous):

yeah thats what is written on my hw

OpenStudy (zzr0ck3r):

the left hand side at x=pi/3 is 2/sqrt{3} the right hand side at x = pi/3 is 4/sqrt(3)

OpenStudy (anonymous):

im confused?

OpenStudy (zzr0ck3r):

im saying if you let x =pi/3 on the left side of the equation we get a different number than doing the same thing on the right side of the equation

OpenStudy (zzr0ck3r):

\[\frac{1-cos(x)}{sin(x)}+\frac{sin(x)}{1-cos(x)}=2csc(x)\] this is true

OpenStudy (anonymous):

must have been a typo

OpenStudy (zzr0ck3r):

\[\frac{1-cos(x)}{1-cos(x)}*\frac{1-cos(x)}{sin(x)}+\frac{sin(x)}{sin(x)}*\frac{sin(x)}{1-cos(x)}\] = \[\frac{(1-cos(x))^2+sin^2(x)}{(1-cos(x))*sin(x)}=\frac{1-2cos(x)+[pcos^2(x)+sin^2(x)]}{(1-cos(x))*sin(x)}\]=\[\frac{1-2cos(x)+[1]}{(1-cos(x))*sin(x)}=\frac{2-2cos(x)}{(1-cos(x))*sin(x)}=\frac{2(1-cos(x))}{(1-cos(x))*sin(x)}=\frac{2}{sin(x)}=\\2csc(x)\]

OpenStudy (zzr0ck3r):

it should say this where it gets cut off \[\frac{2}{sin(x)}=2csc(x)\]

OpenStudy (zzr0ck3r):

let me know if you don't understand anything

OpenStudy (zzr0ck3r):

that pcos(x) should just be cos(x)

OpenStudy (zzr0ck3r):

understand @cgurley86

OpenStudy (anonymous):

not really

OpenStudy (zzr0ck3r):

which part?

OpenStudy (zzr0ck3r):

notice I just multiplied each term by 1?

OpenStudy (anonymous):

no